# Jingyao Liu♠♣  Chen Huang♠♢†  Zhizhao GuanWenqiang Lei♠♣Yang Deng

### v1

@@ Abstract
The rapid advancement in large language models (LLMs) has demonstrated significant potential in End-to-End Software Development (E2ESD). However, existing E2ESD benchmarks are limited by coarse-grained requirement specifications and unreliable evaluation protocols, hindering a true understanding of current framework capabilities. To address these limitations, we present E2EDev, a novel benchmark grounded in the principles of Behavior-Driven Development (BDD), which evaluates the capabilities of E2ESD frameworks by assessing whether the generated software meets user needs through mimicking real user interactions (Figure 1). E2EDev comprises (i) a fine-grained set of user requirements, (ii) multiple BDD test scenarios with corresponding Python step implementations for each requirement, and (iii) a fully automated testing pipeline built on the Behave framework. To ensure its quality while reducing the annotation effort, E2EDev leverages our proposed Human-in-the-Loop Multi-Agent Annotation Framework (HITL-MAA). By evaluating various E2ESD frameworks and LLM backbones with E2EDev, our analysis reveals a persistent struggle to effectively solve these tasks, underscoring the critical need for more effective and cost-efficient E2ESD solutions. Our codebase and benchmark are publicly available at https://github.com/SCUNLP/E2EDev. Refer to caption Figure 1: Overview of E2EDev: a dataset and BDD-based automated evaluation pipeline for E2ESD tasks.
$\dagger$$\dagger$}\$footnotetext: Work done during his PhD program at Sichuan University.
@@ 1 Introduction
The effectiveness of large language models (LLMs) in understanding user needs and demonstrating logical reasoning has been validated in code generation (Jimenez et al., 2024; Gao et al., 2024). LLMs can generate code snippets based on provided code context and user query (Yu et al., 2024; Chen et al., 2021; Hui et al., 2024), and debug code through feedback from either human or system (Zhong et al., 2024; Chen et al., 2023; Li et al., 2022). The success in generating isolated functions is fueling further research into advanced software development automation. This has led to a shift from function-level code synthesis to End-to-End Software Development (E2ESD), where complete software is automatically generated from user requirements (e.g., ‘generate an HTML game of Angry Birds’). By this means, E2ESD could enhance both the code effectiveness and efficiency by removing the manual assembly bottleneck. Current LLM-based E2ESD frameworks can be broadly categorized into multi-agent approaches, inspired by traditional software engineering principles  (Dong et al., 2024; Hong et al., 2024; Du et al., 2024; Royce, 1987; Qian et al., 2024), and single-agent approaches, where a single LLM manages the complete development pipeline (Engineer, 2024; AI, 2024).
As LLM-driven approaches advance, there is growing demand for robust benchmarks to automate the evaluation of E2ESD frameworks (Hong et al., 2024; Qian et al., 2024; He et al., 2024; Hu et al., 2025). However, existing E2ESD benchmarks, such as SoftwareDev (Hong et al., 2024) and SRDD (Qian et al., 2024), suffer from two critical limitations. (1) Coarse-grained Requirement Specifications: Current benchmarks substitute user requirements with ambiguous software descriptions as input, making it difficult to verify whether the generated software aligns with actual user needs. As shown in the top panel of Figure 2, a vague software requirement for Facebook like “sharing with friends” could be interpreted in multiple ways, such as forwarding content, sharing links, or enabling visibility of posts. Without precise operational specifications of such requirements, systematic testing becomes impractical. (2) Unreliable Evaluation Protocols: Evaluations in these benchmarks rely heavily on labor-intensive human assessments (Hong et al., 2024) and lack standardized evaluation methodologies grounded in established software engineering principles (Qian et al., 2024). This results in inconsistent and unreliable performance comparisons across frameworks.
To this end, we introduce the E2EDev benchmark for evaluating the performance of LLM-based frameworks on E2ESD tasks, along with a human-in-the-loop annotation framework designed to ease the annotation burden. Specifically, E2EDev is derived from real-world open-source web application projects. Following the principles of Behavior-Driven Development (BDD), which specifies and validates software behavior from a user perspective, E2EDev evaluates whether the generated software meets user requirements by mimicking real user interactions. As illustrated in the bottom panel of Figure 2, E2EDev consists of: (1) a fine-grained list of user requirements for each software, (2) for each user requirement, we provide multiple BDD test scenarios to verify its correctness, each corresponding to an executable Python code implementation, and (3) an automated testing pipeline built on Behave [^1] Behave is a Python framework for behavior-driven development using natural language tests (Gherkin). See https://github.com/behave/behave for details.. To alleviate the annotation burden while ensuring data quality, we propose the Human-in-the-Loop Multi-Agent Annotation Framework (HITL-MAA). In this framework, specialized agents analyze the project source code to generate candidate requirements and executable tests, with human supervisors involved at key points to avoid errors.
@@
Figure 2: Comparing existing work with ours, existing benchmarks use coarse-grained requirements with unreliable evaluations whereas E2ESD provides fine-grained requirements with executable BDD tests.
Our comprehensive evaluation of various E2ESD frameworks and LLM backbones reveals a significant struggle to effectively solve E2ESD tasks. Current frameworks, even with GPT-4o, exhibit a lack of proficiency in handling detailed functional specifics during implementation, with top performance falling significantly below 60%. Moreover, multi-agent architectures often exhibits excessive interaction rounds and token costs, yielding only minimal gains in effectiveness. Our in-depth error analysis further reveals that the primary challenge lies in accurately implementing functionality with precise details, a challenge amplified using multi-agent architectures. Finally, we believe our benchmark stands out as a valuable resource, highlighting the critical need in designing more effective and cost-efficient E2ESD frameworks.
@@ 2 Related Work
Benchmarks for Software Development. Various benchmarks have been proposed to evaluate the capabilities of LLM-based methods in software development. Most focus on function-level tasks (Chen et al., 2021; Austin et al., 2021; Hendrycks et al., 2021; Zhuo et al., 2024)(upper part of Table 1), while few address project-level E2ESD (lower part). Specifically, SoftwareDev (Hong et al., 2024) relies on human ratings to assess requirement satisfaction across 70 samples, while SRDD (Qian et al., 2024) employs heuristic metrics such as Completeness (absence of unfinished or placeholder code), Executability (compilation and runtime success), and Consistency (semantic similarity between requirements and code). However, these project-level benchmarks face two main limitations. First, the input specifications are often generated by LLMs or presented as high-level textual descriptions (Qian et al., 2024), which are typically vague or underspecified. This makes it difficult to verify whether the generated software truly satisfies the intended requirements. Second, the absence of standardized evaluation methods hinders objective and reliable performance comparison across different frameworks. While rSDE-Bench (Hu et al., 2025) proposes $\sim$600 test cases for project-level coding task evaluation, its reliance on Unit Tests limits assessment to function-level evaluation, overlooking the overall behavior of LLM-generated projects and failing to verify actual user needs. Appendix E details the difference between Unit Tests and our BDD Test approach.
Table 1: Comparison of Representative Benchmarks for LLM-based Software Development
Benchmark Level # Tasks # Reqs With Tests Source Notes
HumanEval (Chen et al., 2021) Function 164 164 Human-Written Python only
MBPP (Austin et al., 2021) Function 974 974 Human-Written Crowd-sourced
APPS (Hendrycks et al., 2021) Function 10k+ 10k+ Real-World Competitive programming
BigCodeBench (Zhuo et al., 2024) Function 1140 1140 Real-World Designed for complex instructions and tool usage
SoftwareDev (Hong et al., 2024) Project 70 - (Human-rated) Real-World Private benchmark
SRDD (Qian et al., 2024) Project 1,200 - (Heuristic) LLM-Generated LLM-generated software description
rSDE-Bench (Hu et al., 2025) Project 53 - (#Unit Tests = 616) LLM-Generated Function-level evaluation with unit tests
E2EDev Project 46 244 (# BDD Test Scenarios = 703) Real-World
Fine-grained user requirements
Project-level evaluation using BDD tests
LLM-based Frameworks for End-to-End Software Development. Recent advances primarily fall into two categories: multi-agent and single-agent frameworks. The majority utilize a multi-agent paradigm, breaking down the development process into subtasks based on established software engineering models (Sommerville, 2011; Hong et al., 2024; Du et al., 2024; Rasheed et al., 2024; Sami et al., 2024). Within this category, Self-Collaboration (Dong et al., 2024) mimics the fundamental development stages—requirement analysis, implementation, and testing—via specialized agents. MapCoder (Islam et al., 2024) enhances software generation by incorporating prior project knowledge. In addition, while most frameworks use predefined workflows, some explore dynamic agent interactions (Qian et al., 2024; Lin et al., 2024), such as ChatDev (Qian et al., 2024), which uses multi-turn dialogues for iterative refinement. Conversely, a smaller body of work pursues a single-agent approach (Engineer, 2024), assuming a single LLM can independently handle the entire development cycle without explicit task decomposition.
@@ 3 E2EDev Benchmark Construction
Max Min Mean Total
Project Task Prompt Length 1907 206 712.74 46
Requirement Reqs per Project 11 2 5.30 244
# Words per Req. 112.76 20 64.10
Test Case Cases per Req. 7 1 2.88 703
Cases per Project 34 2 15.28
[…]
# Lines per Scenario 30 7 11.17
# Words per Scenario 360 53 109.02
[…]
# Lines per Step Definition 157 13 61.41
# Words per Step Definition 700 74 217.44
@@
3.1 Project Source and Selection
To ensure that benchmark data represents realistic and diverse development scenarios while remaining manageable for LLM-based methods, we focus specifically on Web applications as our project source. Web applications strike a balance between complexity and accessibility: they involve rich development aspects such as front-end logic, API communication, and state management, yet typically avoid the need for complex compilation or specialized runtime environments, allowing consistent execution and evaluation (Brown, 2019; Gustafson, 2013; Li et al., 2014).
To ensure the realism and quality of our benchmark, we follow a two-step process for selecting source projects. First, we crawl GitHub repositories using keywords such as “mini project”, “software project”, “web application”, and “H5”, retaining only those with over 500 stars to ensure sufficient community endorsement and visibility. Next, we manually verify each selected project across three key dimensions: (1) executability (runs without major errors in Chrome); (2) functionality (core components behave as intended); and (3) suitability (all core functionalities can be fully experienced within approximately five minutes, thus maintaining manageable complexity). Finally, this comprehensive selection process results in a total of 46 projects. More detailed information regarding the manual verification process can be found in Appendix A.1.
3.2 Human-in-the-Loop Multi-Agent Annotation Framework (HITL-MAA)
@@
User Requirement Annotation. Given the source code of a project, this process involves the collaborative effort of two LLM agents, cooperating with human supervisors: (1) The Code Analyzer Agent analyzes the UI structure and JavaScript logic of the source code to summarize key functionalities and their interactions with UI elements. (2) The Requirement Extractor Agent then automatically generates candidate user-facing requirements based on the summarized analysis. During this process, human supervisors review the generated requirements to ensure accuracy, clarity, and functional consistency through UI inspection or browser testing. In practice, LLM-generated requirements often fail to capture hidden preconditions for software features (e.g., navigation buttons that only appear at smaller screen sizes), which highlights the value of human corrections to prevent the creation of invalid test cases and misleading executable tests later on. Finally, for each data entry in E2EDev, we create an overview summary based on the validated requirements. This provides a description of the project, designed to improve the LLM’s comprehension during our benchmark evaluation.
@@
BDD Test Scenario Annotation. For each human-validated requirement, HITL-MAA generates a corresponding set of Gherkin-style BDD test scenarios, where each scenario describes a specific condition or user interaction to be tested. Scenarios follow the Given-When-Then format, which clearly lays out the preconditions (Given), the user action (When), and the expected outcome (Then) in natural language. To support this process, HITL-MAA employs a Test Case Generation Agent that analyzes the validated requirements and source code to generate relevant scenarios. The agent processes requirements one by one and is prompted to cover a wide range of user behaviors—including typical user actions, unexpected actions, boundary cases, and defensive handling—based on user intent. Since the agent relies on source code, it may miss certain behavior patterns that are not explicitly represented. For example, a function like submitForm() might manage different types of invalid input internally, but if these differences are not clear from the source code, the agent might not generate test scenarios covering every case. To improve coverage, five software testing experts collaboratively review and refine the generated scenarios. They supplement missing interactions, eliminate hallucinated cases, and ensure each scenario is consistent with the requirement. To further ensure annotation quality, HITL-MAA includes a final expert review step to resolve any potential inconsistencies between the requirements and test cases.
Step Implementation Annotation. This step focuses on converting each BDD test scenario into an executable Python script. To achieve this, HITL-MAA utilizes a Test Automation Engineer Agent to generate initial Python step implementations corresponding to the Gherkin-style test steps.
@@
The generated scripts then undergo an iterative verification process to ensure both executability and correctness. Two LLM agents provide feedback to support self-correction: the Dry Run Verifier checks for missing or syntactically invalid step definitions, while the Test Runner executes the script to detect logical inconsistencies or failed assertions. Scripts unresolved after maximum self-correction attempts are flagged for human refinement. Empirically, this self-correction mechanism resolves nearly all dry-run errors autonomously and corrects logical issues in over 80% of cases without human intervention. Validation on the generated dataset achieved a 100% pass rate across all executable tests, demonstrating the high quality and reliability of E2EDev.
3.3 Features & Applications of E2EDev Benchmark
Features of E2EDev. E2EDev consists of diverse software development tasks evenly distributed across multiple common types of web applications and a wide range of common user interaction patterns, detailed in Figure 9 in Appendix A.4. The most prevalent types include: Mathematics & Conversion Tools, such as calculators for tax; Account & Form Management, covering simple CRUD systems where users submit forms and manage records; and Mini Games & Interactive Entertainment like Tic-Tac-Toe. Additionally, E2EDev captures both basic and advanced user interactions, such as clicking, typing, sliders, and drag-and-drop operations. It also includes technically complex features like DOM manipulation, local storage usage, external API integration, and precise mathematical function generation. Taken together, the diversity in application types, interaction patterns, and technical complexity demonstrates that E2EDev provides a representative and realistic benchmark for end-to-end web application development.
Automated Testing Using E2EDev. To enable automated evaluation, we offer a benchmark suite designed for researchers regardless of their familiarity with software testing procedures [^3] Cf. Appendix D for background knowledge of project/software testing..

### v4

# E2EDev: Benchmarking Large Language Models in End-to-End Software Development Task

@@ Abstract
The rapid advancement in large language models (LLMs) has demonstrated significant potential in End-to-End Software Development (E2ESD). However, existing E2ESD benchmarks are limited by coarse-grained requirement specifications and unreliable evaluation protocols, hindering a true understanding of current framework capabilities. To address these limitations, we present E2EDev, a novel benchmark grounded in the principles of Behavior-Driven Development (BDD) to assess whether the generated software meets user needs through mimicking real user interactions. E2EDev comprises (i) a fine-grained set of user requirements for each target software project (ii) multiple BDD test scenarios with corresponding Python step implementations for each requirement, and (iii) a fully automated testing pipeline built on the Behave framework. By evaluating various E2ESD frameworks and LLM backbones with E2EDev, our analysis reveals a persistent struggle to effectively solve these tasks, underscoring the critical need for more effective and cost-efficient E2ESD solutions. Our codebase and benchmark are available at https://github.com/SCUNLP/E2EDev.
E2EDev: Benchmarking Large Language Models in End-to-End Software Development Task
Jingyao Liu1,4  Chen Huang2thanks: Corresponding author.   Zhizhao Guan1,4  Wenqiang Lei1,4  Yang Deng3 1 College of Computer Science, Sichuan University 2 Institute of Data Science, National University of Singapore 3 CHAT NLP Group, Singapore Management University 4 Engineering Research Center of Machine Learning and Industry Intelligence, Ministry of Education, China liujingyao1@stu.scu.edu.cn, huang_chen@nus.edu.sg
@@ 1 Introduction
The effectiveness of large language models (LLMs) in understanding user needs and demonstrating logical reasoning has been validated in code generation (Jimenez et al., 2024; Gao et al., 2024; Zhang et al., 2025a). LLMs can generate code snippets based on provided code context and user query (Yu et al., 2024; Chen et al., 2021; Hui et al., 2024), and debug code through feedback from either human or system (Zhang et al., 2025b; Zhong et al., 2024; Yang et al., 2025; Li et al., 2022; Yang et al., 2024). The success in generating isolated functions is fueling further research into advanced software development automation. This has led to a shift from function-level code synthesis to End-to-End Software Development (E2ESD), where complete software is automatically generated from user requirements (e.g., ‘generate an HTML game of Angry Birds’). By this means, E2ESD could enhance both the code effectiveness and efficiency by removing the manual assembly bottleneck. Current LLM-based E2ESD frameworks can be broadly categorized into multi-agent approaches, inspired by traditional software engineering principles  (Dong et al., 2024; Hong et al., 2024; Du et al., 2024; Royce, 1987; Qian et al., 2024), and single-agent approaches, where one LLM manages the complete development pipeline (Engineer, 2024; AI, 2024).
@@
Figure 1: Comparing existing work with ours, existing benchmarks use coarse-grained requirements with unreliable evaluations whereas E2EDev provides fine-grained requirements with executable BDD tests.
As LLM-driven approaches advance, there is growing demand for robust benchmarks to automate the evaluation of E2ESD frameworks (Hong et al., 2024; Qian et al., 2024; He et al., 2024; Hu et al., 2025). However, existing E2ESD benchmarks, such as SoftwareDev (Hong et al., 2024) and SRDD (Qian et al., 2024), suffer from two critical limitations. (1) Coarse-grained Requirement Specifications: Current benchmarks substitute user requirements with ambiguous software descriptions as input, making it difficult to verify whether the generated software aligns with actual user needs. As shown in Figure 1 (left), the vague requirement “manage words” for a dictionary tool can refer to various actions, such as editing existing entries or bookmarking words. Without precise operational specifications of such requirements, systematic testing becomes impractical. (2) Unreliable Evaluation Protocols: Evaluations in these benchmarks rely heavily on labor-intensive human assessments (Hong et al., 2024) and lack standardized evaluation methodologies grounded in established software engineering principles (Qian et al., 2024). This results in inconsistent and unreliable performance comparisons across frameworks.
To this end, we introduce the E2EDev benchmark for evaluating the performance of LLM-based frameworks on E2ESD tasks, along with a human-in-the-loop annotation framework designed to ease the annotation burden. Specifically, E2EDev is derived from real-world open-source web application projects. Following the principles of Behavior-Driven Development (BDD), which specifies and validates software behavior from a user perspective, E2EDev evaluates whether the generated software meets user requirements by mimicking real user interactions, as illustrated in Figure 1 (right). E2EDev consists of: (1) a fine-grained list of user requirements for each software, (2) for each user requirement, we provide multiple BDD test scenarios to verify its correctness, each corresponding to an executable Python code implementation, and (3) an automated testing pipeline built on Behave [^1] Behave is a Python framework for behavior-driven development using natural language tests (Gherkin). See https://github.com/behave/behave for details.. To alleviate the annotation burden while ensuring data quality, we propose the Human-in-the-Loop Multi-Agent Annotation Framework (HITL-MAA). In this framework, specialized agents analyze the project source code to generate candidate requirements and executable tests, with human supervisors involved at key points to avoid errors.
Our evaluation of various E2ESD frameworks and LLM backbones reveals a significant struggle to effectively solve E2ESD tasks. Even the Claude series, among the strongest coding-oriented LLMs, fails to exceed 60% performance. These failures stem from the difficulty of simultaneously performing high-level planning (i.e., deciding what to build) and faithfully following fine-grained functional details. In multi-agent staged pipelines, this challenge is often addressed by explicitly separating planning from detailed implementation across different phases. However, the effectiveness of such designs critically depends on how intermediate information is communicated. Excessive context sharing tends to dilute key information, whereas overly restricted communication weakens the influence of earlier design choices on later implementation, resulting in high interaction costs with limited performance gains. We believe this benchmark underscores the need for more effective and cost-efficient E2ESD framework design.
@@ 2 Related Work
Benchmark # Tasks # Reqs Test Type / Scale Source Key Features
SoftwareDev (Hong et al., 2024) 70 Amb. Human Eval. Real Private benchmark; human ratings only
SRDD (Qian et al., 2024) 1,200 Amb. Heuristics Gen. Heuristic indicators; no execution
GitTaskBench (Ni et al., 2025) 18 54 Repo Ops. Real Repo operation execution
Mle-Bench (Chan et al., 2024) 72 72 Pipeline Exec. Real Machine Learning pipeline execution
rSDE-Bench (Hu et al., 2025) 53 Amb. 616 Unit Tests Gen. Function-level output checks
E2EDev 46 244 703 BDD Tests Real Requirement Verification via BDD
Table 1: Comparison of Repo-Level Code Generation Benchmarks. Gen.: LLM-Generated. Real: Real-World Repositories. Amb.: Ambiguous or unspecified number of requirements.
Benchmarks for Software Development. Table 1 summarizes representative benchmarks for E2ESD. These benchmarks, which involve project-level code generation, are scarce due to the inherent complexity of their construction, often resulting in limited dataset sizes (Hong et al., 2024; Hu et al., 2025). They typically face two primary limitations. First, vague or underspecified textual specifications (Qian et al., 2024) make it difficult to verify whether the generated software aligns with user requirements. Second, a reliance on subjective human evaluation or heuristic metrics lacks grounding in established software development principles (Hong et al., 2024; Qian et al., 2024), thereby hindering reliable performance assessment. Although some benchmarks incorporate executable evaluation, they are typically confined to specific commands (Ni et al., 2025), workflows (Chan et al., 2024), or individual functions (Hu et al., 2025), failing to verify whether the overall system behavior aligns with the user’s specific intent.
LLM-based Frameworks for E2ESD. Most frameworks employ a multi-agent paradigm, decomposing the development process into explicit subtasks following established software engineering workflows (Sommerville, 2011; Hong et al., 2024; Du et al., 2024; Rasheed et al., 2024; Sami et al., 2024). These workflows can be either predefined (Dong et al., 2024; Islam et al., 2024) or adaptively constructed during interaction (Qian et al., 2024; Lin et al., 2024). In contrast, some work employs single-agent frameworks (Engineer, 2024), where a single LLM manages the entire development process without explicit task decomposition.
@@ 3 E2EDev Benchmark Construction
Max Min Mean Total
Repo. Prompt Len. 1907 206 712.7 46
Req. Reqs / Repo 11 2 5.3 244
Words / Req. 113 20 64.1
Tests Tests / Req. 7 1 2.9 703
Tests / Repo. 34 2 15.3
Gherkin Scenarios
Lines / Scen. 30 7 11.2
Words / Scen. 360 53 109.0
Step Definitions
Lines / Def. 157 13 61.4
Words / Def. 700 74 217.4
@@
3.1 Project Construction
Project Selection. Focusing on Web applications, we selected 46 highly starred GitHub projects that are executable, functionally complete, and suitable for rapid evaluation (details in Appendix A.1).
Features. These projects cover a broad range of Web application types, such as calculators and mini games, and support diverse user interactions, such as typing, clicking, and dragging (cf. Figure 7, Appendix A.4). They also include technically complex features, ranging from simple DOM manipulation to more advanced tasks such as local storage usage and external API integration. Overall, the diversity, technical depth, and realism of these projects make them a qualified and representative ground for evaluating E2E-REP methods.
3.2 HITL-MAA Annotation Framework
@@
User Requirement Annotation. This process transforms a source project into a list of fine-grained user requirements through the collaborative effort of two LLM agents under human supervision. (1) The Code Analyzer Agent analyzes the project’s core functionalities and their interactions with UI elements by reading the source code. (2) The Requirement Extractor Agent then automatically generates candidate user-facing requirements based on this analysis. Human supervisors review the generated requirements to ensure accuracy, clarity, and functional consistency by verifying them against the actual system behavior.
@@
BDD Test Scenario Annotation. For each human-validated requirement, HITL-MAA generates a set of Gherkin-style BDD test scenarios. Each scenario describes a specific condition or user interaction to be tested and follows the Given–When–Then format, clearly specifying the preconditions (Given), the user action (When), and the expected outcome (Then) in natural language. To support this process, HITL-MAA employs a Test Case Generation Agent that analyzes both validated requirements and source code to produce relevant scenarios. The generation process is guided to cover a wide range of test cases, including both typical behaviors and unexpected or edge cases. Since the agent relies solely on source code—which does not explicitly capture all user interaction patterns—some unconventional scenarios may be overlooked and require supplementation by human experts. To ensure comprehensive coverage and high quality, five software testing experts collaboratively review and refine the generated scenarios. Meanwhile, HITL-MAA incorporates a final expert review step to resolve any remaining inconsistencies between requirements and test cases, guaranteeing reliable and high-quality annotations.
@@
Step Implementation Annotation. This step focuses on converting each BDD test scenario into an executable Python script. To achieve this, HITL-MAA utilizes a Test Automation Engineer Agent to generate initial Python implementations corresponding to the Gherkin-style test steps. The generated scripts then undergo an iterative verification process to ensure both executability and correctness. Two agents provide feedback to support self-correction: the Dry Run Verifier checks for missing or syntactically invalid step definitions, while the Test Runner executes the script to detect logical inconsistencies or failed assertions. Scripts unresolved after maximum self-correction attempts are flagged for human refinement. Empirically, this self-correction mechanism resolves nearly all dry-run errors autonomously and corrects logical issues in over 80% of cases without human intervention. Validation on the generated dataset achieved a 100% pass rate across all executable tests, demonstrating the high quality and reliability of E2EDev.
Automated Evaluation. E2EDev adopts the behave framework for fully automated evaluation. Gherkin scenarios are executed through their corresponding Python step implementations to test the generated project against defined requirements, enabling deterministic pass/fail verification of model-generated code. (Details in Appendix D.3.)